# Configurable photonic simulator for quantum field dynamics

### v1

@@ III.2.2 Coherent inputs
Figure 3: For coherent inputs, the OTA reduces to a complex squeezer layer followed by passive linear gates. The adjusted coherent inputs $\alpha^{\prime}$ and the complex-valued squeezing parameters $\zeta=\xi e^{i\Phi}$ are time-dependent, while the beam splitters remain time-independent.
@@ IV Simulating quantum fields
Next, we discuss some representative classes of quantum field theories that the OTA can efficiently simulate. For simplicity, we consider one spatial dimension $d=1$ and periodic boundary conditions. Generalizations to higher dimensions are straightforward: the total mode number scales as $\tilde{N}^{d}$, affecting only the structure of the BS layers. Similarly, boundary conditions are arbitrary.
@@ IV.1 General Hamiltonian structure
It is a central observation that many scalar field Hamiltonians become diagonal in momentum space. In the discretized picture, this is intimately related to the $\phi$-block of the Hamiltonian matrix $H$ corresponding to a circulant (and by definition, real and symmetric) matrix
$\begin{split}H^{\phi}&=\text{circ}(h^{\phi}_{1},h^{\phi}_{2},\dots,h^{\phi}_{N})\\ &=\begin{pmatrix}h^{\phi}_{1}&h^{\phi}_{N}&\dots&h^{\phi}_{3}&h^{\phi}_{2}\\ h^{\phi}_{2}&h^{\phi}_{1}&h^{\phi}_{N}&\dots&h^{\phi}_{3}\\ \vdots&h^{\phi}_{2}&h^{\phi}_{1}&\ddots&\vdots\\ h^{\phi}_{N-1}&\dots&\ddots&\ddots&h^{\phi}_{N}\\ h^{\phi}_{N}&h^{\phi}_{N-1}&\dots&h^{\phi}_{2}&h^{\phi}_{1}\end{pmatrix},\end{split}$ (26)
@@
and typically encode the dispersion of the theory $\lambda_{j}^{\phi}\propto\omega_{j}$.
@@
$\begin{split}(p_{j}^{\text{cos}})_{k}&=\tfrac{2}{\sqrt{N}}\cos\tfrac{2\pi(j-1)(k-1)}{N},\\ (p_{j}^{\text{sin}})_{k}&=\tfrac{2}{\sqrt{N}}\sin\tfrac{2\pi(j-1)(k-1)}{N}.\end{split}$ (28)
For even $N$, there is an additional eigenvalue $\lambda_{N/2+1}=\sum_{k=1}^{N}h^{\phi}_{k}(-1)^{k-1}$ corresponding to the alternating eigenvector $p_{N/2+1}=\frac{1}{\sqrt{N}}\left[1,-1,\dots\right]^{\intercal}$. We emphasize that the eigenvectors are independent of the matrix elements $h_{j}^{\phi}$; hence, the BS layers are identical for all Hamiltonians diagonalized by a DFT.
@@ IV.2 Relativistic theory
with $\epsilon>0$ denoting the lattice spacing and $j\in\mathcal{J}=\{1,\ldots,N\}$ being a discrete label, such that $L=N\epsilon$.
@@ IV.4 Complex fields and the free Bose gas
$\begin{split}\bm{H}_{\text{NR}}=\frac{1}{2}\sum_{j\in\mathcal{J}}\Big{\{}&\frac{1}{2m\epsilon^{2}}\left[(\bm{\phi}_{j+1}-\bm{\phi}_{j})^{2}+(\bm{\pi}_{j+1}-\bm{\pi}_{j})^{2}\right]\\ &+V_{j}\left(\bm{\phi}_{j}^{2}+\bm{\pi}_{j}^{2}\right)\Big{\}}.\end{split}$ (39)
@@
describing ultracold atoms hopping between neighboring sites in an optical lattice, with hopping amplitude $-1/(m\epsilon^{2})$ and a site-dependent potential $2[V_{j}+1/(m\epsilon^{2})]$. The corresponding Hamiltonian matrix decomposes as $H_{\text{NR}}=H^{\phi}_{\text{NR}}\oplus H^{\pi}_{\text{NR}}$, with $H^{\phi}_{\text{NR}}=H^{\pi}_{\text{NR}}$. Since the two blocks agree, we find $\gamma_{j}=1$, such that no squeezing is required. Thus, the non-relativistic theory, or equivalently, the Bose-Hubbard model, can be simulated with passive linear optics only.
@@ IV.5 Curved spacetimes
Since more than two decades, simulating quantum fields in curved spacetime [113, 114] is of substantial interest to experimentally probe phenomena that remain inaccessible in the nightsky, including cosmological particle production [10, 19, 61, 25, 26, 27], the Hawking [9, 56, 12, 21, 115] and Unruh [22] effects, and rotating analog black holes [116, 117]. The standard approach in the field of analog gravity is to consider low-energy perturbations in a fluid, whose effective equation of motion is that of a massless Klein-Gordon field in a curved spacetime dictated by the background fluid flow in the form of the so-called acoustic metric [118]. It is well-known, however, that this one-to-one correspondence breaks down for $1+1$ spacetime dimensions as conformal symmetry prevents the wave equation of the perturbations from being rewritten in terms of a proper metric – thereby restricting many experiments to qualitative analogies. Furthermore, simulating a massive field typically requires two-component fluids [119]. In the following, we demonstrate that a photonic simulation with the OTA overcomes both of these problems.
@@
$\mathcal{L}_{\text{CS}}=\dfrac{1}{2}\sqrt{-g}\left[g^{\mu\nu}(\partial_{\mu}\phi)(\partial_{\nu}\phi)-m^{2}\phi^{2}\right],$ (43)
where $\mu,\nu=\{t,x\}$ label spacetime coordinates (with the Einstein convention understood) and $\sqrt{-g}=\sqrt{-\text{det}(g_{\mu\nu})}$ being the invariant volume element. In two dimensions, every metric is conformally flat and thus proportional to the standard Minkoskwi metric up to a conformal factor, i.e., $g_{\mu\nu}=f^{2}(t,x)\eta_{\mu\nu}=f^{2}(t,x)\text{diag}(1,-1)$. Prominent choices for $f(t,x)$ include static situations, e.g., two-dimensional Anti-de-Sitter space AdS2, for which $f(x)=1/x^{2}$, or, scenarios where a causal horizon arises for the quasi-particle excitations, such as, for instance, $f(x)\propto 1+\tanh(x)$ [21, 115]. On the other hand, there are purely time-dependent settings: choosing an FLRW-type metric $g_{\mu\nu}=\text{diag}[1,-a^{2}(t)]$ encodes cosmological expansion of a homogeneous and isotropic universe [10, 25, 26, 27].
@@
with a spacetime-dependent effective mass $m_{\text{eff}}(t,x)=mf(t,x)$ (for the FLRW metric, we find $f(t)=a(t)$ and the Hamiltonian must be divided by $a(t)$). As was the case for Minkowski spacetime, the Hamiltonian (44) is block diagonal $H_{\text{CS}}=H_{\text{CS}}^{\phi}\oplus H_{\text{CS}}^{\pi}$ with $H_{\text{CS}}^{\pi}=\mathds{1}_{N}/\epsilon$. However, analogous to the spatially varying trapping profile in the non-interacting Bose-Hubbard model [see Eq. (40)], spatial curvature breaks the circulanticity of the $\phi$-block, which is of the form (41) with $\mu_{j}(t)=\epsilon m^{2}_{\text{eff}}(t,\epsilon j)-2\nu$ and $\nu=-1/\epsilon$.
In the absence of spatial curvature, the latter reduces to the relativistic expression (31) with a time-dependent effective mass, which can still be diagonalized by means of the same DFT. More precisely, the fixed form of the DFT eigenvectors (28) guarantees the BS layer to remain unchanged, while the eigenvalues, i.e., the dispersion, changes over time $\omega_{j}=\omega_{j}(t)$, which implies non-linear time-dependencies in the phases $\varphi_{j}(t)$ and the squeezing parameters $z_{j}(t)$. For the FLRW universe, the overall factor $1/a(t)$ can be reabsorbed by mapping the physical time $t$ to $\int_{0}^{t}\mathrm{d}t^{\prime}a(t^{\prime})$, which is reminiscent of the conformal time in cosmology [120]. In static situations, the OTA maintains its form, with all time dependencies isolated in a single phase shift layer, while the circuit parameters must be computed numerically in general.
@@ V Information dynamics
As an application, we study the spreading of entanglement in spacetime after a global quantum quench for varying coupling range. We focus on identifying a regime where continuum features emerge while maintaining the number of modes at an experimentally feasible level.

### v2

@@ III.2.2 Vacuum inputs
Figure 3: For vacuum inputs, the OTA reduces to a complex squeezer layer followed by an interferometer. The complex-valued squeezing parameters $\zeta=\xi e^{i\Phi}$ are time-dependent, while the interferometer remains time-independent. This matches the structure of the Gaussian boson sampler setup.
Thanks to these simplifications, the OTA reduces to a layer of time-dependent squeezers and a single time-independent interferometer array, to wit
$S(t)=S_{\text{I}}S_{\text{CSq}}(\zeta(t)).$ (26)
We sketch the corresponding circuit in Fig. 3, which precisely matches the one realized in Gaussian boson samplers [74, 92]. In such experiments, particle-number-resolving detectors at the output [100, 126, 42] introduce non-Gaussianity. It is well-known that the corresponding photon number output distribution $p(n_{1},\dots,n_{N})=\lvert\braket{n|\bm{U}_{\text{I}}\bm{U}_{\text{CSq}}(t)|0}\rvert^{2}$ contains a Hafnian [74]. Computing Hafnians (and even approximating them) is $\text{P}-$hard in general, with the computational complexity scaling as $\mathcal{O}(n^{3}2^{n})$ in total detected photon number $n$ [20, 75]. Hence, such quantum-optics experiments offer a true quantum advantage over classical algorithms, which has been demonstrated recently [98, 96]. We remark that generalizations of Gaussian boson sampling to coherent inputs have been discussed as well [147]. In this case, Eqs. (25) and (26) continue to hold, but the initial coherent amplitude must be modified, see App. A.4.
While no particular readout is preferred by the OTA, sampling the particle-number distribution experimentally provides access to the full information contained in the (real-space) particle-number sector of the simulated quantum field theory, which complements current approaches based on full-counting statistics [18, 85]. From this, higher-order correlators and information measures can be inferred to probe symmetry-resolution during non-equilibrium dynamics [10, 130]. Therefore, even in the simplest scenario—Gaussian inputs and linear dynamics—the OTA offers a pathway to quantum advantage across a wide range of physically relevant settings.
@@ IV Simulating quantum fields
Next, we discuss some representative classes of quantum field theories that can be simulated using the OTA. For simplicity, we consider one spatial dimension $d=1$ and periodic boundary conditions. Generalizations to higher dimensions are straightforward: the total mode number scales as $\tilde{N}^{d}$, affecting only the interferometer’s structure. Similarly, boundary conditions are arbitrary, see App. D.1 for a discussion of boundary effects.
@@ IV.1 General Hamiltonian structure
A central observation is that many scalar-field Hamiltonians become diagonal in momentum space. In the discretized picture, this is intimately related to the $\phi$-block of the Hamiltonian matrix $H$ corresponding to a circulant (and by definition, real and symmetric) matrix
[…]
@@
and typically encode the dispersion of the theory $\lambda_{j}^{\phi}\propto\omega_{j}$ since often $\lambda_{j}^{\pi}=$ const.
@@
$\begin{split}(p_{j}^{\text{cos}})_{k}&=\sqrt{\tfrac{2}{N}}\cos\tfrac{2\pi(j-1)(k-1)}{N},\\ (p_{j}^{\text{sin}})_{k}&=\sqrt{\tfrac{2}{N}}\sin\tfrac{2\pi(j-1)(k-1)}{N}.\end{split}$ (29)
For even $N$, there is an additional eigenvalue $\lambda_{N/2+1}=\sum_{k=1}^{N}h^{\phi}_{k}(-1)^{k-1}$ corresponding to the alternating eigenvector $p_{N/2+1}=\frac{1}{\sqrt{N}}\left[1,-1,\dots\right]^{\intercal}$. We emphasize that the eigenvectors are independent of the matrix elements $h_{j}^{\phi}$; hence, the interferometer layers are identical for all Hamiltonians diagonalized by a DFT, including time-dependent scenarios.
@@ IV.2 Relativistic theory
with $\epsilon>0$ denoting the lattice spacing and $j\in\mathcal{J}=\{1,\ldots,N\}$ being a discrete label, such that $L=N\epsilon$. We refer to App. D.1.2 for a treatise of open boundary conditions.
@@ IV.4 Complex fields and the free Bose gas
$\begin{split}\hskip-2.84544pt\bm{H}_{\text{NR}}\hskip-1.42271pt=\hskip-1.42271pt\frac{1}{2}\hskip-1.42271pt\sum_{j\in\mathcal{J}}\hskip-1.42271pt\Big\{&\frac{1}{2m\epsilon^{2}}\left[(\bm{\phi}_{j+1}\hskip-1.42271pt-\hskip-1.42271pt\bm{\phi}_{j})^{2}\hskip-1.42271pt+\hskip-1.42271pt(\bm{\pi}_{j+1}\hskip-1.42271pt-\hskip-1.42271pt\bm{\pi}_{j})^{2}\right]\\ &+V_{j}\left(\bm{\phi}_{j}^{2}+\bm{\pi}_{j}^{2}\right)\Big\}.\end{split}$ (40)
@@
describing ultracold atoms hopping between neighboring sites in an optical lattice, with hopping amplitude $1/(m\epsilon^{2})$ and a site-dependent potential $2[V_{j}+1/(m\epsilon^{2})]$. The corresponding Hamiltonian matrix decomposes as $H_{\text{NR}}=H^{\phi}_{\text{NR}}\oplus H^{\pi}_{\text{NR}}$, with $H^{\phi}_{\text{NR}}=H^{\pi}_{\text{NR}}$. Since the two blocks agree, we find $\gamma_{j}=1$, such that no squeezing is required. Thus, the non-relativistic theory, or equivalently, the Bose-Hubbard model, can be simulated with passive linear optics only.
@@ IV.5 Curved spacetimes
Since more than two decades, simulating quantum fields in curved spacetime [19, 105] is of substantial interest to experimentally probe phenomena that remain inaccessible in the nightsky, including cosmological particle production [80, 55, 141, 153, 148, 131], the Hawking [62, 120, 142, 106] and Unruh [79] effects, and rotating analog black holes [143]. The standard approach in the field of analog gravity is to consider low-energy perturbations in a fluid, whose effective equation of motion is that of a massless Klein-Gordon field in a curved spacetime dictated by the background fluid flow in the form of the so-called acoustic metric [15, 29]. It is well known, however, that this one-to-one correspondence comes with several restrictions. First, the acoustic metric has a particular form. Hence, it is not possible to engineer arbitrary spacetimes this way. Second, the mapping breaks down in $1+1$ dimensions, as conformal symmetry prevents the wave equation for the perturbations from being rewritten in terms of a proper metric—thereby limiting the majority of experiments to qualitative analogies [55, 141]. Third, simulating massive fields typically requires two-component fluids [161], which has not yet been demonstrated. In the following, we show that a photonic simulation with the OTA overcomes all three problems.
@@
$\mathcal{L}_{\text{CS}}=\dfrac{1}{2}\sqrt{-g}\left[g^{\mu\nu}(\partial_{\mu}\phi)(\partial_{\nu}\phi)-m^{2}\phi^{2}\right].$ (44)
Here, $x^{\mu}=(t,x)$ labels spacetime coordinates and $\sqrt{-g}=\sqrt{-\text{det}(g_{\mu\nu})}$ renders the volume element invariant under arbitrary coordinate transformations.
A remarkable result from differential geometry is that all two-dimensional manifolds are conformally flat [102]. In general, a pseudo-Riemannian manifold is conformally flat if for each point there exists an open neighborhood that can be mapped to flat, i.e., Minkowski space, via a conformal transformation. These transformations preserve angles but scale lengths locally; the stereographic projection from the $2$-sphere to the plane is a prominent example. Intuitively, conformal flatness amounts to spacetime regions appearing scaled, but never sheared, with respect to flat space. While conformally flat manifolds exist in arbitrary dimensions, two-dimensional manifolds are always conformally flat, which is a consequence of low dimensionality: The metric is symmetric by definition and hence has three degrees of freedom in two dimensions. Two of them single out spacetime coordinates, while the remaining one expresses local rescalings. Hence, the metric is necessarily of the form
$g_{\mu\nu}=f(x^{\mu})\,\eta_{\mu\nu}=f(x^{\mu})\,\text{diag}(1,-1),$ (45)
where $f(x^{\mu})>0$ denotes the conformal factor and $\eta_{\mu\nu}=\text{diag}(1,-1)$ is the Minkowski metric. The high degree of symmetry is also reflected in the Riemann tensor, which quantifies curvature. In two dimensions, it has only one independent component, i.e., $R_{\mu\nu\alpha\beta}=(R/2)(g_{\mu\alpha}g_{\nu\beta}-g_{\mu\beta}g_{\nu\alpha})$, showing that spacetime curvature is entirely determined by the Ricci scalar. Expressed in terms of the conformal factor, the latter reads $R=(1/f)\eta^{\mu\nu}\partial_{\mu}\partial_{\nu}\ln f$.
@@
which is reminiscent of (31) with a spacetime-dependent effective mass $m_{\text{eff}}(x^{\mu})=m\sqrt{f(x^{\mu})}$. In general, this Hamiltonian breaks the circulanticity of the $\phi$-block, which is of the form (42) with $\mu_{j}(t)=\epsilon m^{2}f(t,\epsilon j)-2\nu$ and $\nu=-1/\epsilon$, while $H^{\pi}=\mathds{1}_{N}/\epsilon$.
In what follows, we explore spacetime metrics prominent in general relativity. In particular, we provide coordinate choices favorable for simulation via the OTA, along with the corresponding conformal factors and expressions for the Ricci scalar.
IV.5.1 Variations in time: Cosmology
Let us first consider time-dependent scenarios $f=f(t)$ which are relevant for cosmology. When imposing periodic boundary conditions, the $\phi$-block of (46) boils down to the relativistic result (32) with time-dependent diagonal entries $\mu(t)=\epsilon m^{2}f(t)-2\nu$, thus recovering circulanticity, leading to a now time-dependent relativistic dispersion
$\omega^{2}_{j}(t)=m^{2}f(t)+\dfrac{4}{\epsilon^{2}}\sin^{2}\left[\tfrac{\pi(j-1)}{N}\right].$ (47)
Since the diagonalizing DFT remains time-independent, we can apply the time-dependent OTA (Sec. III.1.3) to simulate cosmological dynamics.
On large scales, our universe is homogenous and isotropic, which is captured by the Friedmann–Lemaître–Robertson–Walker (FLRW) metric $g_{\mu\nu}=\text{diag}[1,-a^{2}(t)]$ [160]. Therein, $a(t)$ specifies the scale factor, which models cosmological expansion and contraction as a function of cosmic time $t$. It is convenient to introduce conformal time $\eta(t)=\int_{0}^{t}\mathrm{d}t^{\prime}/a(t^{\prime})$ [160]. Indeed, when identifying laboratory with conformal time, i.e., considering coordinates $x^{\mu}=(\eta,x)$, the FLRW metric attains its conformally flat form (45) with conformal factor
$f_{\text{FLRW}}(\eta)=a^{2}(\eta).$ (48)
Spacetime curvature manifests in the Ricci scalar as $R=(2/a^{3})a^{\prime\prime}-(2/a^{4})a^{\prime 2}$, where $a^{\prime}(\eta)\equiv\mathrm{d}a(\eta)/\mathrm{d}\eta$.
Hence, when operating the time-dependent OTA in conformal time, cosmological situations with arbitrary scale factors $a(\eta)$ can be engineered. This goes substantially beyond ultracold atom simulations, where only power-law scale-factors have been considered [153, 148]. Cosmologically relevant choices for $a(\eta)$ are summarized in Tab. 1, see also Refs. [19, 105].
IV.5.2 Variations in space: Horizons and (Anti-) de Sitter spaces
Spatial variations break the circulanticity of the $\phi$-block. Yet, time-independent scenarios are always OTA-simulatable, with circuit parameters that must be computed numerically in general. In contrast to cosmology, the coordinates in which the metric assumes its conformally flat form depend strongly on the spacetime of interest; see Tab. 2 for an overview and Refs. [19, 105].
Of particular interest are horizons that causally disconnect a spacetime region from its surroundings. The simplest example is the so-called Rindler spacetime, which describes flat spacetime as perceived by an observer moving at constant acceleration, giving rise to thermal Unruh radiation. Starting from flat-space coordinates $(t,x)$, the accelerated observer uses Rindler tortoise coordinates $(\eta,\rho)$ defined via $t=(e^{\kappa\rho}/\kappa)\sinh(\kappa\eta)$ and $x=(e^{\kappa\rho}/\kappa)\cosh(\kappa\eta)$ covering the right Rindler wedge $x>\lvert t\rvert$.
The prime examples of horizons are black-hole-like metrics of the form $g_{\mu\nu}=\text{diag}[f(r),-1/f(r)]$, where $r$ denotes a radial coordinate and spacetime curvature is given by the simple formula $R=-\mathrm{d}^{2}f(r)/\mathrm{d}r^{2}$. Similar to the Rindler setup, observers near the horizon observe a thermal spectrum—Hawking radiation. Analogously to conformal time, we introduce tortoise coordinates $r^{*}(r)=\int_{0}^{r}\mathrm{d}r^{\prime}/f(r^{\prime})$ to obtain the conformally flat form, in which case $f(r^{*})$ becomes the conformal factor. Note here that one needs to solve the differential equation $r^{*}=r^{*}(r)$ to find $f(r^{*})$. This class encompasses the famous Schwarzschild spacetime $f(r)=1-r_{s}/r$, where $r_{s}=2M$ denotes the Schwarzschild radius in natural units for a black hole of mass $M$. Another prominent variant is the Reissner-Nordström metric for which $f(r)=1-r_{s}/r+(Q/r)^{2}$, describing a black hole of mass $M$ and charge $Q$. We stress that in practice, black-hole-like metrics must feature a cutoff at finite $r>0$ to avoid the singularity at $r=0$.
Type $a(t)$ $a(\eta)$ $R(\eta)$
  Power law $\propto t^{q},q\neq 1$ $\propto\lvert\eta\rvert^{\frac{q}{1-q}}$ $\propto\tfrac{2q}{q-1}(a\eta)^{-2}$
Tunable expansion acceleration: $0\leq q<1$ and $q>1$ describe de- and accelerated expansion, respectively; Includes radiation ($q=1/2$) and matter ($q=2/3$) dominated phases of our universe.
  Milne $\propto t$ $\propto e^{\eta}$ 0
Coasting expansion with vanishing curvature.
  De Sitter $\propto e^{Ht}$ $\propto-(H\eta)^{-1}$ $\propto 2H^{2}$
Describes the rapid inflatory phase of our universe with Hubble rate $H$ and constant positive, i.e., spherical, curvature; Corresponds to the time-like patch of dS2.
  Bouncing $\propto\sqrt{1+t^{2}}$ $\propto\cosh\eta$ $\propto\text{sech}^{4}\eta$
Toy model avoiding the Big Bang singularity as $R(0)=1$ remains finite.
Table 1: Scale factors relevant for cosmology.
Interestingly, space-like (static) patches of (Anti-)de Sitter spacetimes, which correspond to constant curvature, i.e., maximally symmetric, spaces, also fall into the black-hole class: $f(r)\propto 1\pm H^{2}r^{2}$, with (+)-(+) for (Anti-)de Sitter metrics and $H>0$. It is well known that Anti-de Sitter spaces are central to holography [9], underscoring the prospects of the OTA for studying quantum gravity models once gravitational degrees of freedom are included. We remark that there are also alternative coordinate choices. This includes the conformal factor $f(r)\propto 1/(Hr)^{2}$ for AdS2, describing the so-called Poincaré patch, which is conformally flat by construction.
As in cosmological scenarios, the OTA enables analog gravity experiments with spatially dependent metrics far beyond what is achievable with fluid-based approaches [15, 29]. In the latter case, ultracold-atom experiments are restricted to toy models in which horizons arise from interpolating between over- and underdense regions via, for instance, $f(x)\propto 1+\tanh(x)$ [106].
@@ V Information dynamics
As an application, we study the spreading of entanglement in spacetime after a global quantum quench for varying coupling range. We focus on identifying a regime in which continuum features emerge while keeping the number of modes at an experimentally feasible level. Given the classical hardness of probing the particle-number sector, we focus on Gaussian observables.
Type Coordinates $f$ $R$
  Rindler $(\eta,\rho)$ $e^{2\kappa\rho}$ 0
Describes a constantly accelerating observer in flat spacetime experiencing Unruh radiation.
  Black-hole-like $(t,r^{*})$ $f=f(r^{*})$ $\tfrac{\mathrm{d}^{2}f(r)}{\mathrm{d}r^{2}}$
Includes, e.g., Schwarzschild $f(r)=1-r_{s}/r$ and Reissner-Nordström $f(r)=1-r_{s}/r+(Q/r)^{2}$ solutions featuring event horizons and Hawking radiation.
  De Sitter $(t,r^{*})$ $\propto\text{sech}^{2}(Hr^{*})$ $\propto 2H^{2}$
Static patch of a maximally symmetric space with constant positive curvature and $H>0$.
  Anti-De Sitter $(t,r^{*})$ $\propto\text{sec}^{2}(Hr^{*})$ $\propto-2H^{2}$
Static patch of a maximally symmetric space with constant negative curvature; Relevant for holographic dualities.
Table 2: Conformal factors encoding spatially-varying metrics.